OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4. Medal?

OpenStudy (anonymous):

@aum Hey I just ran into another question I need help with. Can you help?

OpenStudy (aum):

First distribute the -2 on the left hand side as well as the right hand side. Tell me what you get?

OpenStudy (anonymous):

x+3=x+1-4?

OpenStudy (anonymous):

I'm not sure what you mean by distribute. I always get confused with the word distribute

OpenStudy (aum):

No. The -2 outside means each term inside the parenthesis has to be multiplied by -2. That is what distribute -2 means here.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

-2x-3=-2x-1-4

OpenStudy (aum):

Not -3. -2 * 3 = -6 Left hand side is: -2(x + 3) = -2x - 6 Right hand side is: -2(x + 1) - 4 = -2x - 2 - 4 = -2x - 6 Equate the left and right hand sides: -2x - 6 = -2x - 6 Both LHS and RHS are identical. This tell ys that no matter what x is, the LHS and RHS will ALWAYS be equal. That means this equation has infinite number of solutions. You can put any value for x and the given equation -2(x + 3) = -2(x + 1) - 4 will ALWAYS be true. Therefore there are infinite solutions where x can be ANY real number.

OpenStudy (anonymous):

So all real numbers?

OpenStudy (aum):

Yes x can be ANY real number.

OpenStudy (anonymous):

Oh. Ok, thank you for the help again!

OpenStudy (aum):

you are welcome.